OpenStudy (anonymous):

PLEASES HELP NOW!! Add -4 3/8 + -5 3/4= -10 1/8 -8 1/8 -1 3/8 -3/8

OpenStudy (texaschic101):

-4 3/8 - 5 3/4 I am going to set aside the whole numbers and deal with the fractions... -3/8 - 3/4 = -3/8 - 6/8 = -9/8 = -1 1/8 now the whole numbers...-5 +(-4) = -9 now add them, whole numbers + fractions -9 + (-1 1/8) = -10 1/8

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you help me with another

OpenStudy (anonymous):

@texaschic101